The tabular environment with horizontal and vertical lines bounding left-justified cells with vertical lines (|l|l|) and horizontal lines (\hline):

\begin{tabular}{|l|l|l|}
\hline
No. & Name & Function \\
\hline
1.  & CONTACT & to define a contact \\
2.  & END     & to end the data input \\
\hline
\end{tabular}

No. Name  Function
1. CONTACT to define a contact
2. END  to end the data input

The tabular environment again, this time with prescribed column widths:

\begin{tabular}{p{1.2in}p{4.0in}}
NAME    & Required string \\
        & A string giving the name of the contact used in the. \\
        & solution module.\\
SURFACES & Required stringlist \\
         & A stringlist giving the names of the surfaces and
windows which define the contact.  It is not necessary for both
contacts and windows to be used in the definition.
\end{tabular}

NAME  Required string

A string giving the name of the contact used.

SURFACES  Required stringlist

A stringlist giving the names of the surfaces and windows which define the contact. It is not necessary for both contacts and windows to be used in the definition.